Consider a monopolistic competitor who has the following demand function P=140-4Q. In addition suppose that his total cost function is given by TC=10+5Q2.
Required:
i) Find his equilibrium price (P) and quantity (Q). ​​​​
ii) What is his maximum profit? ​​​​​
b) Explain the features of monopolistic completion.​​​​]
REVISION QUESTION
a) A perfectly competitive firm is faced with the following total cost schedule
Q​= 0​1​2​3​4​5​6​7​8​9​10​

TC​= 9​20​30​39​47​54​60​67​77​90​109

Required:

(i) If the market price is $13, what output will the firm choose to produce to maximize profit? What is the maximum profit?
1
Expert's answer
2020-04-16T09:52:16-0400

a) MR"=" MC;

MC"=10Q;"

"MR=((140-4Q)\u00d7Q )'=(140Q-4Q^2)'=140-8Q;"

"10Q=140-8Q; \n18Q=140;"

"Q=" "7.8;"

"P=140-8\u00d77.8; P=140-31.2;"

"P=" $"108.8;"

Profit"=TR-TC;"

"TR=7.8\u00d7108.8=" $ "848.64;"

"TC=10+5\u00d77.8^2=" $ "314.2;"

"Profit=848.64-314.2=" $ "534.44";